# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ 4.1.3. Parsing Agent
The agent’s primary responsibility is to translate step descriptions into a sequence of formalized instructions tailored for smartphone operations. Building on previous work (Li et al., 2020a), the definition of instruction has been enriched to include a more detailed object description. As shown in Table 4.1.3, an instruction is represented as a tuple comprising the operation type, associated parameters, and object descriptions, usually referencing a widget on the page. The agent utilizes LLM for parsing, because of its capacity to function effectively with minimal initial data and its adeptness in handling flexible text types (e.g., natural language, several keywords). As the number of cases increases, the instruction definition can be extended and the agent can be progressively refined (further discussed in Section 4.2.3). Users can validate the parsed results and, where necessary, can perform manual corrections.
@@ 4.1.4. Grounding Agent
The final stage aims to predict specific operations on a smartphone based on provided instructions. Specifically, it identifies a UI widget on the smartphone with the object description in the instruction and thus executes the relevant operation.
Initially, the agent captures the current page’s layout hierarchy and screenshot. Based on text and interactivity attributes, it filters potential candidate widgets. Subsequently, to ensure an efficient and effective prediction of widgets, the agent adopts a multi-tiered strategy:
Historical RPA invocation: If the provided instructions are recurring, the agent consults the historical RPA repository (in Section 4.2.1) to identify widgets used in the past. It then compares these widgets with the ones on the current page, using heuristic rules that focus on their rich GUI attributes. A widget is considered a match only if it exceeds a certain level of similarity.
New instruction-based grounding: For new instructions, the agent’s sole reliance is on the object description in the instruction. The agent then employs the following strategies:
@@
    Textual semantic matching: The agent identifies potential widgets, focusing on semantic similarity with the operation and object description in the instruction. It leverages LLM to predict the widget most suited to execute the given instruction.
@@
    Exploratory mechanism: Challenges in identifying the operation widget can emerge due to mismatches between the actual page and the provided instructions, often a result of outdated or erroneous instructions, variable initial states of smartphone screens, or unexpected changes on the page (Figure 4). Consequently, the agent does not strictly adhere to the instructions for executing operations. Instead, our exploratory mechanism leverages a wealth of contextual information to guide the decision-making process of the LLM, enabling it to infer the most probable current operation. Furthermore, the mechanism allows the agent to make several trials on the smartphone, reducing total reliance on provided instructions. Issues such as compound instructions (Figure 7) and inconsistent starting points (Figure 7) can be effectively addressed. As the agent persistently accrues knowledge and insights, the contextual information becomes richer, solidifying the foundation upon which the LLM bases its inferences (further discussed in Section 4.2.4).
Figure 7. Two examples illustrating the effectiveness of the exploratory mechanism. (a) The compound instruction Click “Sign up” requires additional instructions. The agent can supplement the necessary steps. (b) The initial state doesn’t match the first instruction. The agent can skip unnecessary instructions.
For the implementation details of each strategy, see Appendix A.1. After the agent makes its widget prediction, users have the option to manually demonstrate the correct operation on the mobile GUI if they find the prediction to be inaccurate.
4.1.5. Mobile Semantics Agent
The agent complements the Grounding Agent by enhancing its understanding of mobile semantics, a critical factor for successful grounding. There is a discernible heterogeneity between text-based instructions and the operations on the mobile page. GUI contains structured information and images but often lacks textual alignment with instructions, which is most evident when widgets, particularly icons, are void of textual descriptions.
To bridge this divide, the agent enriches the mobile page semantically, using a combination of layout hierarchy and screenshots to create textual descriptions for the page and its widgets. We employ the screen2words model (Wang et al., 2021) to generate full-page textual descriptions and a combination of OCR and pix2struct (Lee et al., 2023) for widgets. Based on these models, the agent is fine-tuned through continuous absorption of new training data (further discussed in Section 4.2.1).
@@ 4.1.6. Assessment Agent
The agent is introduced to proactively address potential pitfalls and streamline decision-making. It ensures the Grounding Agent remains focused on its primary tasks without being overloaded with error detection. The agent adopts three key strategies:
  • Evaluation of execution: It assesses whether operations executed up to the present are reasonable, to estimate the success of the ongoing task. The proactive assessment provides the Grounding Agent with informed guidance on whether to continue or retrace steps.
  • Redundancy Prevention: By marking historical attempts, the agent avoids falling into repetitive cycles, thereby preventing endless loops during the exploration.
  • Intelligent User Involvement: It gauges the confidence in current decisions to determine whether user intervention is appropriate. Figure 8 illustrates a scenario where the agent prompts user intervention due to decision uncertainty. Similarly, Figure 1 showcases a situation where the instruction “Click a ringtone you like” introduces an ambiguity that necessitates user assistance.
The agent leverages the LLM for reasoning. Based on the contextual information, such as the current page description, operation history, and immediate instructions, the LLM decides whether to proceed with the current operation or roll back, providing a confidence score for its decision. More details are in Appendix A.2.
@@ 4.2. Accumulated Knowledge
The agents draw from an ever-expanding knowledge base, which undergoes refinement as user usage and intervention accumulate. As a feedback mechanism, the knowledge base not only boosts the agents’ learning but also enhances the entire system’s efficiency and task success rates.
Based on the guidelines in Section 3.4, the four categories of knowledge integral to PromptRPA are as follows:
  • Historical RPA Repository stores successfully executed RPA tasks. The validated tasks not only demonstrate a higher success rate in future runs but also retain adjustments tailored to specific operating environments or update previously outdated tutorials, minimizing the need for manual reconfigurations.
@@
  • Instruction Set focuses on improving the system’s ability to extract instructive information from text. Failed parsing examples become valuable learning materials.
@@
We will introduce them in detail and outline strategies for the continuous improvement they bring to the system.
4.2.1. Historical RPA Repository
Upon successful execution, PromptRPA archives the RPA task model’s data, according to the structure defined in Section 3.1. Given that the Grounding Agent may not always align operations with initial instructions, the repository focuses on storing the final sequence of operations. These sequences are then reverse-engineered into corresponding instructions and step descriptions via simple mapping rules, as detailed in Appendix B.1.
The historical RPA repository yields several optimizations across the system:
  • The Retrieval Agent will prioritize historical RPA task models, enhancing data reliability and minimizing retrieval time from online sources.
  • The Grounding Agent gains enhanced predictive capabilities. It can utilize a richer feature set, incorporating layout hierarchies and image data alongside text descriptions. Notably, the stored mapping between instructions and operations acts as a learning resource for the agent, facilitating the fine-tuning of its internal model. This repository-driven learning equips the Grounding Agent to more accurately identify widgets when similar operation descriptions are encountered in the future.
  • The Mobile Semantics Agent benefits from a richer set of textual descriptions for pages and widgets, using them to fine-tune its models. The agent archives images and their embeddings (Wang et al., 2021) from the repository. By comparing new images with archived ones, the agent improves its accuracy in generating textual descriptions for similar images encountered later.
@@ 4.2.2. Context Library
    User-Driven Updates: The library continuously updates based on user-system interactions. For instance, the earphones mentioned by the user in a certain use (Figure 10) will be collected into the library.
In the library, a unique set of parameter records is maintained for each user. The Analysis Agent starts by identifying the contextual parameters required for the specific task. It then consults the library, prioritizing the user’s own historical records over those of other users. If required parameters are missing, the user is prompted to provide the missing details, ensuring the task’s function description is completed.
Figure 9. An example of records in the context library.
Figure 10. An example of auto-completing the textual prompt.
Figure 10 showcases a sample set of these parameter records, which will be dynamically updated based on user interaction. With the aid of the library, as depicted in Figure 10, missing parameters for recurring textual prompts are automatically supplemented, which users have the option to confirm.
@@ 4.2.3. Instruction Set
The Parsing Agent is tasked with converting natural language step descriptions into formalized instructions. While conceptually straightforward, practical application introduces issues, including the emergence of new operation, parameter, or object description types and synonyms or deviation from the format in Table 4.1.3. To address these issues, the instruction set is regularly updated to include misinterpreted cases and their corrections. These cases can come from user manual corrections, but most are resolved by instructions reverse-engineered from successful RPA tasks, as described in Section 4.2.1. Continuous refinement of the instruction set aids in the ongoing fine-tuning of the model within the Parsing Agent.
Another important issue is related to compound instructions (Figure 7). For such cases, the instruction set records both the compound instruction and its decomposed result. Hence, if a compound instruction reoccurs, the Parsing Agent will automatically decompose it.
@@ 4.2.4. Mobile Interaction Graph
In the graph, each node represents a visited mobile page, and edges link interactive widgets to the pages they lead to (Deka et al., 2017). As the system navigates more pages, the graph evolves dynamically. The graph enables virtual validation of each instruction before its real-world execution on a smartphone. If the edge already exists in the graph, it anticipates the future pages that could be visited after the operation, thus enriching the contextual understanding for the Grounding Agent and the Assessment Agent, facilitating more informed inferences (see Appendix A for details).
@@ 4.3. Implementation
The system is divided into two primary components: a server and a client. The server handles information collection, instruction generation, and accumulated knowledge. It is built using the Flask framework and crawls search results from Baidu, Bing, and Google by simulating web browser actions. On the client installed on the smartphone, the Accessibility Service API  [^2] https://developer.android.com/reference/ android/accessibilityservice/AccessibilityService is used to obtain page layout data, and MediaProjection  [^3] https://developer.android.com/reference/android/media/projection/MediaProjection is used to capture screenshots. The client triggers operations through the Accessibility Service API on the smartphone. When user intervention is required, an overlay is provided to monitor the user’s demonstrations. Actions captured by this overlay are transferred to the native mobile GUI to trigger the corresponding operations. The PromptRPA’s agents use the GPT-4 API  [^4] https://platform.openai.com/docs/models/gpt-4 for their language model needs and GPT-3.5 Turbo model for fine-tuned needs. All involved prompts are shown in the supplemental files.
5. Pipeline Performance Evaluation
This section assesses the performance of PromptRPA’s three-stage pipeline—information collection, instruction generation, and operation mapping—as well as the overall system performance.
@@ 5.1. Experiment Setup
PromptRPA is designed to fulfill a wide range of real-world needs through textual prompts provided by users, which involve various applications and built-in functionalities. Therefore, we specifically created a dataset for the experiment.
Targeting Chinese speakers, we selected 10 Chinese applications (9 applications and 1 system) for evaluation, as shown in Table 2. Five of them are top-ranked in terms of active users as of 2021, and four have rankings beyond 300. The diverse selection enables us to evaluate PromptRPA’s performance across different user activity levels and application complexities. We constructed 100 ground-truth RPA task models by the following steps:
  • We collected 37,987 questions related to these applications by crawling several Q&A websites. The questions were then ranked by their click-through rates, and we selected 5 common and 5 uncommon questions for each application, resulting in a total of 100 questions. Each question was used to construct one RPA task model and each was manually annotated with a ground-truth function description.
@@
  • We utilized crowdsourcing to collect textual prompts that participants would use to generate RPA tasks. In total, we collected 25 different textual prompts for each RPA task, amounting to 2,500 prompts. Among them, 100 prompts contained complete step descriptions, negating the need for further tutorial retrieval.
  • According to the step description, we then manually annotated the ground-truth instructions for each RPA task model.
  • Finally, the experiment was conducted on a HUAWEI Mate30. For each RPA task model, we manually recorded the ground-truth operation sequence on this device. It is crucial to note that these sequences are not annotated based on ground-truth instructions, which might be impractical due to issues like outdated tutorials.
Table 2. List of applications used and examples of tasks in the experiment. WeChat, QQ, TikTok, Beautiful Weather, Safe and Sound, and Film Encyclopedia were also selected for the user evaluation in Section 6.
Application Category Activity Ranking Task Examples
  WeChat Message 1 1. View WeChat Bean details

### v2

@@ 4.1.3. Parsing Agent
The agent’s responsibility is to translate step descriptions into a sequence of formalized instructions tailored for smartphone operations. Building on previous work (Li et al., 2020a), the definition of instruction has been enriched to include a more detailed object description. As shown in Table 4.1.3, an instruction is represented as a tuple comprising the operation type, associated parameters, and object descriptions, usually referencing a widget on the page. The agent utilizes LLM for parsing, because of its capacity to function effectively with minimal initial data and its adeptness in handling flexible text types (e.g., natural language, several keywords). As the number of cases increases, the instruction definition can be extended and the agent can be progressively refined (further discussed in Section 4.2.3). As shown in Figure 11, users can validate the parsed results and, where necessary, can perform manual corrections.
@@
Figure 8. Operation mapping workflow. The Grounding Agent predicts the operation and the Assessment Agent performs the final check. The colors of the components within the prompt indicate their sources, which are the same color elements pointing to the prompt with the same color arrows. Gray arrows represent the normal task flow, while purple arrows indicate the accumulation of knowledge.
@@ 4.1.4. Grounding Agent
As shown in Figure 8, the final stage aims to predict specific operations on the smartphone given the instructions. Specifically, according to the object description in the instruction, it identifies the UI widget on the smartphone and thus executes the relevant operation.
Initially, the agent filters out all widgets that lack both interactive attributes and text descriptions. Interactive attributes include being clickable, long clickable, editable, focusable, checkable, and scrollable. Text descriptions include both explicit text and implicit content descriptions. They are all available in the layout hierarchy. Then the agent relies on the mobile semantics module to bridge the gap between text-based instructions and GUI described by layout hierarchy and screenshots. Considering real-time efficiency, the module employs the screen2words model (Wang et al., 2021) to generate full-page textual descriptions and utilizes a combination of OCR and pix2struct (Lee et al., 2023) to describe widgets which lack text descriptions.
Subsequently, to predict the operation efficiently and effectively, the agent adopts a multi-layer strategy as follows:
Historical task invocation: If the given instructions are recurring, the agent consults the historical task repository (in Section 4.2.1) to identify widgets used in the past. It then compares these widgets with the ones on the current page, using heuristic rules that compare their rich GUI attributes. A widget is considered a match only if it exceeds the preset similarity thresholds.
New instruction-based grounding: For new instructions or when the historical task invocation fails, the agent employs the following strategies:
@@
    If this strategy fails, the agent then employs the LLM to decide which of the remaining five strategies to use: redirect, add, skip, expand, and block.
@@
    Redirect: The agent leverages LLM to predict the widget most suited to execute the given instruction. It focuses on semantic similarity with the operation and object description in the instruction.
    However, instructions and operations do not always correspond one to one. To make the mapping more flexible, the agent can choose from the following exploratory strategies:
@@
    Add: The agent incorporates additional steps when necessary, often due to outdated or erroneous instructions, or when an instruction step requires expansion (Figure 9). The agent leverages as much contextual information as possible to infer the most probable operation rather than solely relying on the initial instructions. When the agent predicts an incorrect operation, the Assessment Agent can help it retract and potentially find the correct operation. This strategy is crucial when there are unexpected changes on the page (Figure 3).
  4. (4)
    Skip: The agent skips unnecessary or irrelevant steps to better align the current page state with the instructions. This strategy helps in addressing inconsistent starting points, allowing the process to continue smoothly even if some steps are not applicable (Figure 9).
  5. (5)
    Expand: When all instructions have been finished, or the remaining instructions are totally useless, the agent keeps trying to explore potential operations, until the Assessment Agent indicates task completion.
  6. (6)
    Block: The agent is currently unable to proceed as it cannot identify a suitable interactive widget, often due to the lack of information on the current page. In such cases, the agent requests user intervention by describing the current situation and the expected user actions. The user is asked to manually demonstrate the correct operation on the mobile GUI, as shown in Figure 11. If the user chooses to ignore, the agent will re-invoke the LLM to employ one of the above strategies.
Figure 9. (a) The compound instruction Click “Sign up” requires additional instructions. The agent can add the necessary steps. (b) The state doesn’t match the first instruction. The agent can skip unnecessary instructions.
As the agent persistently accrues knowledge and insights, it improves mobile semantics extraction and LLM reasoning (further discussed in Section 4.2.4). For the implementation details of each strategy, see Appendix A.4.
@@ 4.1.5. Assessment Agent
The agent is introduced to proactively address potential pitfalls and simplify decision-making. It ensures the Grounding Agent remains focused on its primary tasks without being overloaded with error detection.
As shown in Figure 8, for efficiency, historical task invocations are thought to be sufficiently reliable and thus do not require assessment, allowing for direct execution. For new instruction-based grounding, the agent first assesses whether the current operation is a previously failed attempt, preventing endless loops during exploration. Subsequently, the agent evaluates whether the operations executed up to the present are reasonable. The agent leverages the LLM for reasoning. Based on contextual information including the current page description, operation history, and the current instruction, the agent adopts one of the following strategies:
  • Follow: The agent determines that the operation is reasonable and proceeds with execution on the smartphone.
  • Change: The agent determines that the operation does not meet the current task requirements and instructs the Grounding Agent to remove the erroneous option and make a new prediction.
  • Retract: If the Grounding Agent has previously attempted erroneous expansions leading to irrelevant pages, the agent performs a global back operation to return to the previous step.
  • Finish: Usually, tasks are not completed simply by executing all instructions, due to outdated tutorials or errors in preceding stages. Therefore, the Grounding Agent will automatically choose the “expand” strategy to identify further necessary operations after all instructions have been executed. Based on the contextual information, the Assessment Agent determines whether the task is actually completed. The execution process only terminates when the Assessment Agent confirms task completion.
However, sometimes the agent is not sufficiently confident in adopting any of the above strategies. In such cases, we consider it appropriate to seek user assistance. Therefore, the agent employs the LLM to assess if it is confident in its decision. If the agent is unconfident, it describes its current decision and prompts the user to demonstrate the correct operation. Figure 10 illustrates a scenario where the agent prompts user intervention due to low confidence for decision. Similarly, Figure 1 showcases a situation where the instruction “Click a ringtone you like” introduces an ambiguity that necessitates user assistance. More details are in Appendix A.5.
@@
Notably, even if the agent prompts to seek user assistance, the user has the option to ignore the prompt, in which case the agent will directly execute according to its own decision.
@@ 4.2. Accumulated Knowledge
As shown in Figure 5, 7, 8, the agents draw from an accumulated knowledge base, which is continually improved through user usage and intervention. As a feedback mechanism, the knowledge base facilitates the learning of the agents, thereby enhancing the entire system’s efficiency and task success rates.
According to the above agents’ knowledge requirements, Prompt2Task requires four types of knowledge:
  • Historical Task Repository stores successfully executed automation tasks, which typically exhibit a higher success rate in future runs. Additionally, it retains adjustments tailored to specific operating environments or updates to previously outdated tutorials, eliminating the need for manual reconfigurations upon reuse.
@@
  • Instruction Set focuses on improving the system’s ability to extract instructive information from text. Similar examples become valuable learning materials.
@@
We will introduce them in details and explain how they bring improvements to the agents.
4.2.1. Historical Task Repository
Upon successful execution, Prompt2Task archives the Prompt2Task model’s data, according to the definition in Section 3.1. Since the Grounding Agent may not always align operations with the original instructions, the repository uses the sequence of operations as the ground truth. These operations are then reverse-engineered into corresponding instructions and step descriptions via simple mapping rules, as detailed in Appendix B.1.
The historical task repository provides several optimizations to the agents:
  • The Retrieval Agent will prioritize historical task models, enhancing data reliability and avoiding the time on retrieving online tutorials whenever possible. The agent identifies the most similar historical task models based on the current function description. Only when the similarity exceeds the preset empirical threshold is the model considered a match and directly used for subsequent processing.
  • The Grounding Agent gains enhanced predictive capabilities. For instructions stored in the repository, it can now identify corresponding UI widgets through a richer feature set that incorporates layout hierarchies and screenshots, rather than relying solely on text descriptions from instructions (as exemplified in Case 1 in Appendix B.3). As long as the features of the UI widgets do not change significantly, the heuristic rules employed by the agent can accurately and quickly identify the widgets and trigger the appropriate operations. Additionally, the stored mapping between instructions and operations serves as a learning resource for the agent. If the historical task invocation fails, the agent finds the historical instruction most similar to the current one and then incorporates the corresponding mapping into the prompt, thereby aiding the LLM in decision-making (for details, refer to the prompt engineering of the Grounding Agent in Appendix A.4).
  • The mobile semantics module in the Grounding Agent benefits from a richer set of textual descriptions for widgets. For the widgets’ screenshots stored in the repository, the module computes their embeddings (Wang et al., 2021). When new images with similar embeddings to the stored ones appear, the module uses the stored textual descriptions for them, which are more accurate. It is particularly effective for recognizing icons; as long as the icons’ images maintain a similarity that does not exceed an empirical threshold, the agent can accurately label them when they are encountered again (as exemplified in Case 2 in Appendix B.3).
@@ 4.2.2. Context Library
    User-Driven Updates: The library continuously updates based on user-system interactions, primarily involving parameters mentioned by users in historical tasks, such as connected devices or Wi-Fi names.
In the library, a unique set of parameter records is maintained for each user (refer to the example in Appendix B.2). To support user-driven updates, after each completed automation task, the original records and the current task model are provided to the LLM to generate the latest records.
The Analysis Agent accesses the library to find the user’s historical records and incorporates these records as part of the prompt (for details, refer to the prompt engineering of the Analysis Agent in Appendix A.1). With the aid of the library, missing parameters for recurring textual prompts are automatically supplemented, which users have the option to confirm. For instance, the earphones mentioned by the user in a certain use (refer to Case 3 in Appendix B.3) will be collected into the library and automatically suggested to the user in future tasks. Additionally, if the user’s context changes and the parameters saved in the context library become outdated, users can correct them while chatting with the Analysis Agent, and the parameters in the context library will be updated after the completion of the new task.
@@ 4.2.3. Instruction Set
The Parsing Agent is tasked with converting natural language step descriptions into formalized instructions. While conceptually straightforward, practical application introduces issues, including the emergence of new operation, parameter, or object types and synonyms or deviation from the format in Table 4.1.3. To address these issues, the instruction set is regularly updated to include misinterpreted cases and their corrections. These cases can be resolved by user manual corrections, but most are resolved by instructions reverse-engineered from successful automation tasks, as described in Section 4.2.1. Similar to the historical task repository, the instruction set stores the mapping between the description of a single step and the instruction it is parsed into. For each step that needs parsing, the agent identifies the most similar description from the set, and incorporates the historical mappings into the prompt for the LLM, thereby enhancing the LLM’s parsing capabilities.
Another important issue is related to compound instructions (as exemplified in Figure 9 and Case 4 in Appendix B.3). For such cases, the user marks the compound instruction and its decomposed result, which are then recorded in the instruction set. Hence, if a compound instruction reoccurs, the Parsing Agent will automatically decompose it.
@@ 4.2.4. Mobile Interaction Graph
In the graph, each node represents a visited mobile page, and edges link interactive widgets to the pages they lead to (Deka et al., 2017). As the system navigates more pages, the graph evolves dynamically. Specifically, for a completed automation task, the operation sequence naturally corresponds to a series of page transitions. To represent these transitions in the graph, Prompt2Task uses a heuristic rule-based similarity function (Huang et al., 2023) that considers the texts and their locations on pages, to calculate their similarity and categorize the pages. The similarity thresholds have been established using Receiver Operating Characteristic (ROC) curve analysis. Similar pages share the same node in the graph, and edges between these nodes represent the operation objects that trigger transitions between pages.
The graph enables virtual validation of each operation before actually executing on a smartphone. If the edge already exists in the graph, it anticipates the future pages that could be visited after the operation, thus enriching the contextual understanding for the Grounding Agent and the Assessment Agent (see Appendix A.4 and A.5 for prompt details). As long as the page transition relationships remain unchanged, the graph can provide the agents with accurate information about future pages (as exemplified in Case 5 in Appendix B.3).
@@ 4.3. Implementation
Figure 11. Runtime screenshots of the Prompt2Task client. (a) The user chats with the Analysis Agent to finalize the function description. (b) The Retrieval Agent provides 5 tutorials for the user to choose from. (c) The Parsing Agent generates instructions, which the user can edit. In (a) (b) (c), if the user prefers not to intervene, they can click “START” to skip. (d) The Grounding Agent fails to find the target widget and requests the user to provide a demonstration. The user can begin the demonstration by selecting “CONFIRM”. (e) The Assessment Agent is unconfident in the current predicted operation and requests user intervention, but the user can still choose “IGNORE”, allowing the agent to proceed with its prediction.
Prompt2Task is divided into two primary components: a server and a client. The server handles information collection, instruction generation, and knowledge accumulation. It is developed using the Flask framework and crawls search results from Baidu, Bing, and Google by simulating web browser actions. The client is an Android application installed on smartphones, designed to interact with users and execute automation tasks. As illustrated in Figure 11, the client provides a GUI for user input, selection, modification, and demonstration. During task execution, it runs in the background on the smartphone and utilizes the Accessibility Service API  [^3] https://developer.android.com/reference/android/accessibilityservice/AccessibilityService to access page layout data, and the MediaProjection API  [^4] https://developer.android.com/reference/android/media/projection/MediaProjection to capture screenshots. The client triggers operations on the smartphone through the Accessibility Service API. When user demonstration is required, an overlay is provided to monitor the user’s demonstrations. Actions captured by this overlay are recorded by the system and then transferred to the native mobile GUI to trigger the corresponding operations. Prompt2Task’s agents utilize the GPT-4  [^5] https://platform.openai.com/docs/models/gpt-4 API for their large language model needs. All involved prompts are detailed in the appendix. In tasks involving textual similarity, including the Retrieval Agent finding the most relevant historical task model, the Parsing Agent finding the most similar step descriptions, and the Grounding Agent finding the most similar instructions, we use LaBSE (Feng et al., 2022), which is a state-of-the-art multilingual sentence embedding model, offers an effective trade-off between efficiency and accuracy.
5. Performance Evaluation
This section assesses the performance of Prompt2Task’s three-stage pipeline—information collection, instruction generation, and operation mapping—as well as the overall system performance.
@@ 5.1. Experiment Setup
Prompt2Task is designed to fulfill a wide range of real-world needs through textual prompts provided by users, which involve various applications and built-in functionalities. Therefore, we specifically created a dataset for the experiment.
Targeting Chinese speakers, we selected 10 Chinese applications (9 applications and 1 system) for evaluation, as shown in Table 3. Five of them are top-ranked in terms of active users as of 2021, and four have rankings beyond 100. The diverse selection enables us to evaluate Prompt2Task’s performance across different user activity levels and application complexities. We constructed 100 ground-truth Prompt2Task models by the following steps:
  • We collected 37,987 questions related to these applications by crawling a recognized online Q&A platform [^6] https://jingyan.baidu.com/. The questions were sorted based on their view counts. The top half were categorized as common, while the bottom half were categorized as uncommon. We manually selected 5 common and 5 uncommon questions for each application, ensuring all $10\times 10$ questions were unique. Each question was used to construct one Prompt2Task model and each was manually annotated with a ground-truth function description.
@@
  • We utilized crowdsourcing to collect textual prompts that participants would use to generate automation tasks. The tasks were hosted on a public crowdsourcing website we developed, where participants were required to formulate prompts based on the ground-truth function descriptions provided to them. Each participant was instructed to ensure the uniqueness and relevance of their submissions. After filtering out duplicate, low-quality, or irrelevant prompts, we collected 25 different textual prompts for each automation task, for a total of $25\times 100$ prompts. Among them, 100 prompts contained complete step descriptions, without the need for further tutorial retrieval.
  • According to the step description, we then manually annotated the ground-truth instructions for each Prompt2Task model, totaling 533 instructions.
  • Finally, the experiment was conducted on an HONOR V20. For each Prompt2Task model, we manually recorded the ground-truth operation sequence on this device, totaling 543 operations. Notably, these operations do not directly correspond to the ground-truth instructions, often due to outdated tutorials. Details are available in our public dataset.
Table 3. Applications and tasks overview in the experiment. WeChat, QQ, TikTok, Beautiful Weather, Safe and Sound, and Film Encyclopedia were also selected for the user evaluation in Section 6.
Application Category Activity Step Task Examples
Ranking Range
  WeChat Message 1 4~10 1. View WeChat Bean details